OpenStudy (liliakarina):

(2*5^2)^5/5^10*2^3 I say 10^10/10^13

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

\[\left( 2*5^{2}\right)^{5}\div5^{10}*2^{3}\] thats the problem right?

OpenStudy (liliakarina):

Yes, that is

OpenStudy (anonymous):

ok solve the parenthesis first.

OpenStudy (anonymous):

so it eould be\left( 2*25 \right)\]\[^{5}

OpenStudy (anonymous):

\[50^{5}\]

OpenStudy (anonymous):

pemdas. exponets is the e. so you do \[5^{2}\] not \[\left( 2*5 \right)^{2}\]

OpenStudy (liliakarina):

Ah, okay

OpenStudy (anonymous):

sorry if ui confused you. but it the 5 first and from there you multiply that 2 so its 50 to the fifth power divided by the rest of the prob

OpenStudy (liliakarina):

So, its 50^5/10^13

OpenStudy (anonymous):

ok now on the bottom you have\[5^{10}*2^{3}\]

OpenStudy (anonymous):

naw not that.

OpenStudy (anonymous):

ok now the \[2^{3}\] equals 8 sonow the bottom is \[5^{10}*8\]

OpenStudy (anonymous):

so now the prob is \[50^{5}\div5^{10}*8\]

OpenStudy (liliakarina):

Right. Then would it be 40^10 ?

OpenStudy (anonymous):

ok you cant multipy bases.

OpenStudy (liliakarina):

So it cant get any more simplified?

OpenStudy (anonymous):

\[2*3^{3}\]is six cubed but \[2*3^{3}\] is difeferent cuz it would be 2*9=18 but back to the prob. it can... there a shortcut in to doing it without solvibng all exponets but i dont remember it.

OpenStudy (liliakarina):

Dont worry. You did help me!

OpenStudy (anonymous):

haha ok